# MaskRIS: Semantic Distortion-aware Data Augmentation for Referring Image Segmentation

### v1

@@ 1 Introduction
Referring Image Segmentation (RIS) [16] involves precisely delineating objects within an image based on text descriptions. Unlike semantic and instance segmentation, which are constrained to pre-defined classes, RIS offers unique flexibility by segmenting objects specified by free-form expressions. This capability has led to extensive applications in diverse domains, such as language-driven human-robot interaction [43] and advanced image editing [3, 39]. One of the main challenges in RIS is effectively bridging the modality gap between visual and language features. Recent research [33, 56, 49] has focused on developing sophisticated architectures for cross-modal alignment.
@@
While significant progress has been made in aligning different modalities, effective training techniques for RIS remain underexplored. The trend toward developing larger and more complex models requires challenging training settings like data augmentation. Previous studies [4, 5, 47] have demonstrated the crucial role of data augmentation in model training. However, designing data augmentation for RIS is challenging: standard augmentation used in semantic segmentation is often incompatible with RIS, as some textual descriptions directly conflict with augmentation.
As illustrated in Fig. 1, spatial augmentation, such as random crop and horizontal flip, alters the meaning of “left”, causing the target mask to become incorrect after augmentation. Similarly, color distortion also conflicts with color-specific descriptions like “red”. Due to the incompatibilities, naïvely applying augmentations is not beneficial - our empirical analysis reveals that conventional data augmentation degrades RIS [33] performance (Fig. 2(a)). This finding explains why previous RISs [22, 49, 56, 33, 52, 61] have taken a cautious approach to augmentation, predominantly relying on simple resizing while avoiding complex techniques that could severely distort the data.
@@
To fully exploit the advantages of input masking, we propose a Distortion-aware Contextual Learning (DCL) framework that processes inputs through two complementary paths: a primary path with original inputs and a secondary path with masked inputs. The primary path ensures baseline training stability and maintains model accuracy, while the secondary path enhances data diversity and robustness. The distillation loss aligns predictions from both paths, encouraging the model to make consistent predictions on both original and masked inputs. This DCL framework enhances feature robustness to masked inputs through the secondary path, while the primary path provides stability and mitigates potential harmful effects of severe distortion. Additionally, DCL acts as a regularizer, adding beneficial challenges to the RIS task. The regularization effect aligns with findings from [38], where self-distillation is shown to constrain model representations, reducing overfitting and enhancing robustness - a process similarly achieved by aligning original and masked predictions in our framework.
Our contributions are as follows: (1) We provide a comprehensive analysis of data augmentation techniques for robust RIS training, identifying limitations in previous methods for the first time. (2) Building on these insights, we introduce Masked Referring Image Segmentation (MaskRIS), a simple yet effective baseline that overcomes semantic conflicts of conventional augmentations while increasing data diversity. This framework maximizes the benefits of original and masked inputs for stable training and enhanced accuracy. (3) Our MaskRIS achieves significant performance gains over existing RIS methods, including weakly supervised approaches, setting new state-of-the-art results on RefCOCO, RefCOCO+, and RefCOCOg datasets.
@@ 2 Related Work
Referring Image Segmentation. Referring Image Segmentation (RIS) aims to segment objects in images based on text descriptions. This field evolved from early CNN-RNN/LSTM combinations for image and text processing [37, 45, 16, 26] to advanced approaches using cross-modal attention [29, 32, 57, 17, 18, 9] and pre-trained transformer encoders [49, 22, 56, 33]. While previous studies have focused on image-text alignment, our study shifts focus to the training strategy in RIS, offering a complementary approach to enhance existing studies. This orthogonal perspective allows us to enjoy the performance benefits established by previous methods. Our concurrent work, NeMo [10], introduces a mosaic-based augmentation that combines a target image with carefully selected negative samples. While effective, it requires additional overhead for constructing and retrieving from a negative sample pool. In contrast, our approach provides a more efficient framework that enhances data diversity with a broader analysis of augmentation types and model robustness.
@@
Masked Input Training. Masked input training is a self-supervised learning technique where random parts of input data are masked, and the model learns to predict these masked elements from visible data. Originally successful in natural language processing (NLP) as masked language modeling (MLM) [7, 34], it expanded to computer vision as masked image modeling (MIM) showing both training efficiency and strong generalization performance [2, 53, 11]. These masking strategies are also employed to develop robust feature representations. RandomErasing [62] and Cutout [8] erase randomly selected rectangle areas to reduce over-fitting and handle occlusion. MIC [15] leverages context clues from the target domain through masked image consistency for domain adaptation. SMKD [27] uses masked knowledge distillation for few-shot classification.
@@
Self-distillation. Self-distillation leverages the model itself as both teacher and student, preventing the need for separate teacher networks to facilitate knowledge distillation [14]. PS-KD [21] demonstrates the utility of softening hard targets through teacher model predictions, acting as an effective regularization strategy. Some studies [60, 40] explored self-distillation by employing an entire network as the teacher and an early exit network as the student. CoSub [48] notably improved model performance in visual recognition tasks with sub-model-based self-distillation. MaskSub [13] introduced a drop-based technique for sub-model self-distillation, improving model performance and cost efficiency for image classification tasks. To fully exploit the benefits of input masking, we employ self-distillation between the predictions of the original and masked input.
(a) Performance with conventional augmentations
(b) Performance with individual augmentation on CARIS [33]
@@ 3.1 Referring Image Segmentation (RIS)
RIS Pipeline. Given an image $x\in\mathbb{R}^{H\times W\times 3}$ and a text description that specifies an object within the image, the RIS model outputs a pixel-level mask delineating the referred object. The text description is first tokenized into a sequence of words $w=\{w_{1},w_{2},\ldots,w_{L}\}$, where $w_{i}\in\mathbb{R}^{D}$ represents a word embedding with $D$ as the embedding dimension and $L$ as the number of words in the description. Here, $H$ and $W$ denote the height and width of the image, and $3$ represents the RGB color channels. We formalize this process as follows:
@@
In conventional RIS approaches, $f$ consists of the image encoder [35, 41] and the text encoder [7, 41] to extract high-dimensional visual and language features, along with a vision-language fusion module that integrates these features to generate the segmentation mask. Training the RIS models is essentially a pixel-wise binary classification task using the cross-entropy loss function:
@@
where $N$ denotes the total number of pixels, $y_{i}$ is the binary label at each pixel $i$ in the ground truth mask (1 indicating the object, 0 the background), and $\hat{y}_{i}$ is the predicted probability of the pixel belonging to the object.
Limited Data Augmentation in RIS. Previous studies [56, 49, 33] have only focused on aligning vision and language features to bridge the modality gap. Still, effective training techniques for RIS remain underexplored, limiting the potential of vision-language features. As vision-language models grow in size and complexity [41, 19, 1], optimizing these training techniques becomes more crucial. Among training techniques, we regard data augmentation as the most promising approach, with the potential to enhance model robustness and generalization [42, 24, 59, 4, 5].
Current RIS methods [49, 54, 56, 33] employ only simple resizing operations for data augmentations and don’t utilize complex augmentations used in semantic segmentation. This is due to the fact that the complex data augmentations in segmentation are incompatible with RIS as shown in Fig. 1, often altering or eliminating the referred objects. Consequently, these augmentations lead to a decline in RIS model performance on RefCOCO as demonstrated in Fig. 2. Thus, we believe tackling the limited data augmentation problem is essential for overcoming a major performance bottleneck in RIS.
@@ 3.2 Masking Strategy for Image and Text
To overcome the limitations, we introduce a masking-based augmentation that combines both image and text masking to generate diverse image-text training pairs. Unlike conventional data augmentation, which often distorts essential spatial and attribute details, image masking minimizes such distortions, preserving this critical information while significantly expanding data diversity. In addition, image masking is beneficial in recognizing partially visible or occluded objects, as illustrated in Fig. 3(a) and (b). State-of-the-art models, such as CARIS [33], often struggle to identify partially visible objects. This issue highlights a prevalent challenge: the presence of occlusions often leads to the failure of models trained on clear and unoccluded objects. Similarly, as seen from Fig. 3(b), text descriptions often rely on contextual objects (e.g., “spoon handle” to describe the position of “broccoli”), suggesting that the impact of occlusion extends beyond the target objects themselves.
Figure 3: The existing RIS method tends to be inaccurate when faced with occluded context. CARIS [33] represents the SoTA method in RIS. Words highlighted in red represent occluded objects in the image (left and center) and masked words in the text query (right).
@@
Beyond managing image occlusions, RIS requires flexibility in interpreting varied text descriptions, as a single referred object can be described in multiple ways. This diversity in textual description is essential for generalizing to different referring expressions. While our image masking addresses visual occlusions, we introduce text masking to enhance the model’s capability to interpret diverse and incomplete descriptions by training it to rely on broader contextual cues. By masking parts of the text, we encourage the model to infer missing details, improving its understanding and reducing dependence on specific terms. For instance, as shown in Fig. 3(c), CARIS can accurately identify a “middle guy light blue tie” with complete text. However, it struggles with partial descriptions, such as “[MASK] guy light blue tie”, indicating the limitations in handling incomplete textual clues. By incorporating text masking, we enable the model to adapt to various referring expressions, enhancing its comprehension and robustness in RIS tasks.
@@
We next describe our image and text masking techniques in detail, demonstrating how they address the aforementioned challenges in RIS.
@@
Image Masking. We divide the input image $x$ into non-overlapping patches and randomly mask a fixed ratio of patches. Following the MAE sampling strategy [11], we create a binary mask $M\in\mathbb{R}^{H\times W}$ by sampling patches uniformly without replacement, where 1 indicates a masked patch. The masked image $x^{M}$ is obtained by element-wise multiplication: $x^{M}=(1-M)\odot x,$ where $\odot$ denotes the element-wise multiplication operation. This mechanism ensures that only a subset of the image’s patches is exposed to the model during training, compelling the neural network to infer missing information and thereby enhancing its robustness and predictive accuracy.
@@
Text Masking. We mask word tokens to generate masked word tokens $w^{M}=\{w^{M}_{1},w^{M}_{2},\ldots,w^{M}_{L}\}$ using a probabilistic strategy in MLM [7]. Specifically, each word $w_{i}$ has a 15% chance of being masked. Of these, 80% are replaced with a [MASK] token, making it invisible to the model during training; 10% are replaced with a random word from the vocabulary, introducing noise and variability; and 10% remain unchanged, preserving the original context. This strategy allows the model to experience a variety of textual scenarios, enhancing its ability to comprehend context and meaning even when parts of the text are masked or altered. Notably, as Wei and Zou [50] witnessed, random word masking effectively preserves sentence meaning, further validating our method’s balance between sentence coherence and data diversity.
@@ 3.3 Distortion-aware Contextual Learning
To further facilitate the RIS training, we introduce Distortion-aware Contextual Learning (DCL), as shown in Fig. 4. This approach uses a primary path for processing original inputs and a secondary path for masked inputs. The primary path focuses on preserving the original image context, ensuring training stability, and promoting distortion-aware training. Meanwhile, the secondary path introduces variability by processing masked data, which improves the model’s robustness. The training objective of MaskRIS consists of two parts, the distillation loss $\mathcal{L}_{dist}$ and the original pixel-wise classification loss $\mathcal{L}_{ce}$ in Eq. 2, i.e., $\mathcal{L}_{total}=\mathcal{L}_{dist}+\mathcal{L}_{ce}.$
@@
where $\mathsf{sg}$ is the stop-gradient function, which blocks gradient back-propagation. Inspired by [13], MaskRIS employs a self-distillation framework that leverages soft targets from the original inputs, reducing training complexity and enhancing model generalization. This approach allows MaskRIS to effectively mitigate the learning instability caused by input masking while maintaining the benefits of increased data diversity and robustness.

### v2

@@ 1 Introduction
Referring Image Segmentation (RIS) (hu2016segmentation) involves precisely delineating objects within an image based on text descriptions. Unlike semantic and instance segmentation, which are constrained to pre-defined classes, RIS offers unique flexibility by segmenting objects specified by free-form expressions. This capability has led to extensive applications in diverse domains, such as language-driven human-robot interaction (Robotics) and advanced image editing (ImageEditing1; ImageEditing2). One of the main challenges in RIS is effectively bridging the modality gap between visual and language features. Recent research (CARIS; LAVT; CRIS) has focused on developing sophisticated architectures for cross-modal alignment.
@@
While significant progress has been made in aligning different modalities, effective training techniques for RIS remain underexplored. The trend toward developing larger and more complex models requires challenging training settings like data augmentation. Previous studies (autoaug; randaug; touvron2021training) have demonstrated the crucial role of data augmentation in model training. However, designing data augmentation for RIS is challenging: standard augmentation used in semantic segmentation is often incompatible with RIS, as some textual descriptions directly conflict with augmentation.
As illustrated in Figure 1, spatial augmentation, such as random crop and horizontal flip, alters the meaning of “left”, causing the target mask to become incorrect after augmentation. Similarly, color distortion also conflicts with color-specific descriptions like “red”. Due to the incompatibilities, naïvely applying augmentations is not beneficial - our empirical analysis reveals that conventional data augmentation degrades RIS (CARIS) performance (Figure 3(a)). This finding explains why previous RIS studies (ReSTR; CRIS; LAVT; CARIS; Group-RES; CoupAlign) have taken a cautious approach to augmentation, predominantly relying on simple resizing while avoiding complex techniques that could severely distort the data.
@@
To fully exploit the advantages of input masking, we propose a Distortion-aware Contextual Learning (DCL) framework that processes inputs through two complementary paths: a primary path with original inputs and a secondary path with masked inputs. The primary path ensures baseline training stability and maintains model accuracy, while the secondary path enhances data diversity and robustness. The distillation loss aligns predictions from both paths, encouraging the model to make consistent predictions on both original and masked inputs. This DCL framework enhances feature robustness to masked inputs through the secondary path, while the primary path provides stability and mitigates potential harmful effects of severe distortion. Additionally, DCL acts as a regularizer, adding beneficial challenges to the RIS task. The regularization effect aligns with findings from mobahi2020self, where self-distillation is shown to constrain model representations, reducing overfitting and enhancing robustness - a process similarly achieved by aligning original and masked predictions in our framework.
Figure 2: Comparison with CM-MaskSD (wang2024cm). (a) CM-MaskSD employs global feature guided masking via a vision-language model (VLM), such as CLIP (CLIP), which requires a pre-aligned vision–language representation. It focuses on high-correlation regions but lacks diversity and architecture-agnostic properties. (b) MaskRIS adopts holistic context masking, yielding richer diversity and improved architecture independence.
Finally, we note that although MaskRIS shares certain surface-level similarities with the concurrent work CM-MaskSD (wang2024cm), it is conceptually and technically distinct. As illustrated in Figure 2, CM-MaskSD employs masking guided by cross-modal correlations, where a global text feature is compared against all image patch tokens (and vice versa for a global image feature with text tokens). This correlation-based masking selects the most text-relevant patches and distills them through the [CLS]-token representation, resulting in an object-centric, token-level self-distillation scheme. However, this design inherently relies on a pre-aligned vision–language representation, limiting its generalizability. In contrast, MaskRIS introduces holistic context masking across both image and text, which not only preserves spatial cues and avoids augmentation-induced semantic flips (e.g., “left” and “right”, color changes), but also yields richer diversity by exposing the model to a broader range of masked contexts. Our Distortion-aware Contextual Learning (DCL) aligns segmentation outputs from a clean primary path and a masked secondary path, enforcing pixel-wise consistency while maintaining stability and diversity. This enables reasoning under incomplete or ambiguous multi-modal cues; empirically, masking both modalities outperforms masking only one (see Sec. 4.3). Importantly, MaskRIS is free from pre-aligned cross-modal features, architecture-agnostic, and adds no parameters or inference overhead, thereby departing clearly from prior masking-based RIS methods.
Our contributions are as follows: (1) We provide a comprehensive analysis of data augmentation techniques for robust RIS training, identifying limitations in previous methods for the first time. (2) We propose a holistic context masking strategy that perturbs both image and text, increasing data diversity through varied, yet coherent partial contexts while reducing semantic conflicts across modalities. (3) We introduce Distortion-aware Contextual Learning (DCL), a dual-path distillation framework that enforces pixel-level consistency between original and masked predictions. Unlike prior [CLS] token–dependent schemes tied to specific architectures, our approach combines holistic masking with DCL to enable stable and architecture-agnostic training. (4) Our method achieves significant performance gains over existing RIS methods, including weakly supervised approaches, establishing new state-of-the-art results on RefCOCO, RefCOCO+, and RefCOCOg datasets. Moreover, we show that the proposed framework also benefits Referring Expression Comprehension (REC), highlighting its broader applicability beyond RIS.
@@ 2 Related Work
Referring Image Segmentation. Referring Image Segmentation (RIS) aims to segment objects in images based on text descriptions. This field evolved from early CNN-RNN/LSTM combinations for image and text processing (margffoy2018dynamic; shi2018key; hu2016segmentation; li2018referring) to advanced approaches using cross-modal attention (liu2017recurrent; liu2021cross; ye2019cross; hu2020bi; hui2020linguistic; ding2021vision) and pre-trained transformer encoders (CRIS; ReSTR; LAVT; CARIS). While previous studies have focused on image-text alignment, our study shifts focus to the training strategy in RIS, offering a complementary approach to enhance existing studies. This orthogonal perspective allows us to enjoy the performance benefits established by previous methods. Our concurrent work, NeMo (NeMo), introduces a mosaic-based augmentation that combines a target image with carefully selected negative samples. While effective, it requires additional overhead for constructing and retrieving from a negative sample pool. In contrast, our approach provides a more efficient framework that enhances data diversity with a broader analysis of augmentation types and model robustness.
@@
Masked Input Training. Masked input training is a self-supervised learning technique where random parts of input data are masked, and the model learns to predict these masked elements from visible data. Originally successful in natural language processing (NLP) as masked language modeling (MLM) (BERT; ROBERTA), it expanded to computer vision as masked image modeling (MIM) showing both training efficiency and strong generalization performance (BEIT; SimMIM; MAE). These masking strategies are also employed to develop robust feature representations. RandomErasing (RandomErasing) and Cutout (CUTOUT) erase randomly selected rectangle areas to reduce over-fitting and handle occlusion. MIC (MIC) leverages context clues from the target domain through masked image consistency for domain adaptation. SMKD (SMKD) uses masked knowledge distillation for few-shot classification.
Recently, MagNet (chng2024mask) applies masked text training to improve vision-language alignment in RIS, where masked words are predicted using image features. Similarly, CM-MaskSD (wang2024cm) applies masked input training using object-centric masking and CLIP-based patch-text alignment to refine feature representations. Although both methods improve cross-modal learning, their main focus is on feature alignment. In contrast, MaskRIS is inspired by the semantic conflicts from data augmentation in RIS. To address this issue, we performed a systematic analysis of augmentation-induced inconsistencies and their impact on RIS performance. Based on these findings, MaskRIS introduces a holistic masking strategy and a distortion-aware training framework, explicitly designed to improve model robustness in RIS.
@@
Self-distillation. Self-distillation leverages the model itself as both teacher and student, preventing the need for separate teacher networks to facilitate knowledge distillation (KD). PS-KD (PSKD) demonstrates the utility of softening hard targets through teacher model predictions, acting as an effective regularization strategy. Some studies (zhang2019your; phuong2019distillation) explored self-distillation by employing an entire network as the teacher and an early exit network as the student. CoSub (CoSub) improved model performance in visual recognition tasks with sub-model-based self-distillation. MaskSub (MaskSub) introduced a drop-based technique for sub-model self-distillation, improving model performance and cost efficiency for image classification tasks. To fully exploit the benefits of input masking, we employ self-distillation between the predictions of the original and masked input.
(a) Performance w/ conventional augmentations
(b) Performance w/ individual augmentation on CARIS
@@ 3 Method
In this section, we propose a novel training strategy for RIS, called Masked Referring Image Segmentation (MaskRIS). MaskRIS aims to address the limitation of previous RIS (Sec. 3.1) with two key components: (1) input masking strategy (Sec. 3.2) for data augmentation in RIS and (2) a dual-path training (Sec. 3.3) to maximize the benefits of the masking strategy.
@@ 3.1 Referring Image Segmentation (RIS)
RIS Pipeline. Given an image $x\in\mathbb{R}^{H\times W\times 3}$ and a text description specifying an object within the image, the RIS model outputs a pixel-level mask delineating the referred object. The text description is first tokenized into a sequence of words $w=\{w_{1},w_{2},\ldots,w_{L}\}$, where $w_{i}\in\mathbb{R}^{D}$ represents a word embedding with $D$ as the embedding dimension and $L$ as the number of words in the description. Here, $H$ and $W$ denote the height and width of the image, and $3$ represents the RGB channels. We formalize this process as follows:
@@
In conventional RIS approaches, $f$ consists of the image encoder (Swin; CLIP) and the text encoder (BERT; CLIP) to extract high-dimensional visual and language features, along with a vision-language fusion module that integrates these features to generate the segmentation mask. Training the RIS models is essentially a pixel-wise binary classification task using the cross-entropy loss function:
@@
where $N$ denotes the total number of pixels, $y_{i}$ is the binary label at each pixel $i$ in the ground truth mask (1 for object, 0 for background), and $\hat{y}_{i}$ is the predicted probability of the pixel belonging to the object.
Limited Data Augmentation in RIS. Previous studies (LAVT; CRIS; CARIS) have only focused on aligning vision and language features to bridge the modality gap. Still, effective training techniques for RIS remain underexplored, limiting the potential of vision-language features. As vision-language models grow in size and complexity (CLIP; ALIGN; FLAMINGO), optimizing these training techniques becomes more crucial. Among training techniques, we regard data augmentation as the most promising approach, with the potential to enhance model robustness and generalization (rebuffi2021data; li2023data; yun2019cutmix; autoaug; randaug).
Current RIS methods (CRIS; ETRIS; LAVT; CARIS) employ only simple resizing operations for data augmentations and don’t utilize complex augmentations used in semantic segmentation. This is due to the fact that the complex data augmentations in segmentation are incompatible with RIS as shown in Figure 1, often altering or eliminating the referred objects. Consequently, these augmentations lead to a decline in RIS model performance on RefCOCO as demonstrated in Figure 3. Thus, we believe tackling the limited data augmentation problem is essential for overcoming a major performance bottleneck in RIS.
Figure 4: The existing RIS method tends to be inaccurate when faced with occluded context. CARIS (CARIS) represents the SoTA method in RIS. Words highlighted in red represent occluded objects in the image (left and center) and masked words in the text query (right).
@@ 3.2 Masking Strategy for Image and Text
To overcome the limitations, we introduce a masking-based augmentation that combines both image and text masking to generate diverse image-text training pairs. Unlike conventional data augmentation, which often distorts essential spatial and attribute details, image masking minimizes such distortions, preserving this critical information while significantly expanding data diversity. In addition, image masking is beneficial in recognizing partially visible or occluded objects, as illustrated in Figure 4(a) and (b). State-of-the-art models, such as CARIS (CARIS), often struggle to identify partially visible objects. This issue highlights a prevalent challenge: the presence of occlusions often leads to the failure of models trained on clear and unoccluded objects. Similarly, as seen from Figure 4(b), text descriptions often rely on contextual objects (e.g., “spoon handle” to describe the position of “broccoli”), suggesting that the impact of occlusion extends beyond the target objects themselves.
@@
Beyond managing image occlusions, RIS requires flexibility in interpreting varied text descriptions, as a single referred object can be described in multiple ways. This diversity in textual description is essential for generalizing to different referring expressions. While our image masking addresses visual occlusions, we introduce text masking to enhance the model’s capability to interpret diverse and incomplete descriptions by training it to rely on broader contextual cues. By masking parts of the text, we encourage the model to infer missing details, improving its understanding and reducing dependence on specific terms. For instance, as shown in Figure 4(c), CARIS can accurately identify a “middle guy light blue tie” with complete text. However, it struggles with partial descriptions, such as “[MASK] guy light blue tie”, indicating the limitations in handling incomplete textual clues. By incorporating text masking, we enable the model to adapt to various referring expressions, enhancing its comprehension and robustness in RIS tasks.
@@
We next describe our image and text masking techniques in detail, demonstrating how they address the aforementioned challenges in RIS. In the following, we specify the detailed process for both image and text masking, clarifying how these techniques are implemented to overcome the identified challenges and enhance the RIS model’s performance.
@@
Image Masking. We divide the input image $x$ into non-overlapping patches and randomly mask a fixed ratio of patches. Following the MAE sampling strategy (MAE), we create a binary mask $M\in\mathbb{R}^{H\times W}$ by sampling patches uniformly without replacement, where 1 indicates a masked patch. The masked image $x^{M}$ is obtained by element-wise multiplication: $x^{M}=(1-M)\odot x,$ where $\odot$ denotes the element-wise multiplication operation. This mechanism ensures that only a subset of the image’s patches is exposed to the model during training, compelling the neural network to infer missing information and thereby enhancing its robustness and predictive accuracy.
@@
Text Masking. We mask word tokens to generate masked word tokens $w^{M}=\{w^{M}_{1},w^{M}_{2},\ldots,w^{M}_{L}\}$ using a probabilistic strategy in MLM (BERT). Specifically, each word $w_{i}$ has a 15% chance of being masked. Of these, 80% are replaced with a [MASK] token, making it invisible to the model during training; 10% are replaced with a random word from the vocabulary, introducing noise and variability; and 10% remain unchanged, preserving the original context. This strategy allows the model to experience a variety of textual scenarios, enhancing its ability to comprehend context and meaning even when parts of the text are masked or altered. Notably, as EDA witnessed, random word masking effectively preserves sentence meaning, further validating our method’s balance between sentence coherence and data diversity.
@@ 3.3 Distortion-aware Contextual Learning
To further facilitate the RIS training, we introduce Distortion-aware Contextual Learning (DCL), as shown in Figure 5. This approach uses a primary path for processing original inputs and a secondary path for masked inputs. The primary path focuses on preserving the original image context, ensuring training stability, and promoting distortion-aware training. Meanwhile, the secondary path introduces variability by processing masked data, which improves the model’s robustness. The training objective of MaskRIS consists of two parts, the distillation loss $\mathcal{L}_{dist}$ and the original pixel-wise classification loss $\mathcal{L}_{ce}$ in Eq. 2, i.e., $\mathcal{L}_{total}=\mathcal{L}_{dist}+\mathcal{L}_{ce}.$ In practice, we adopt a weighting factor between $\mathcal{L}_{ce}$ and $\mathcal{L}_{dist}$. While our default setting uses $\lambda=0.5$, we also provide a detailed sensitivity analysis of this ratio in Appendix A.5, showing that MaskRIS is robust across a wide range of choices.
@@
where $\mathsf{sg}$ is the stop-gradient function, which blocks gradient back-propagation. Inspired by MaskSub, MaskRIS employs a self-distillation framework that leverages soft targets from the original inputs, reducing training complexity and enhancing model generalization. This approach allows MaskRIS to effectively mitigate the learning instability caused by input masking while maintaining the benefits of increased data diversity and robustness.